Caffeine in beverages has been a hot topic for some time. It takes time for the body to process caffeine (assume first-order kinetics). The half-life of caffeine in the body is 4.90 hours. Caffeine is also found in chocolate. If a person eats a chocolate bar with 30.0 mg of caffeine at the beginning of a 3.00 hour test, how much (in mg) would be left in their system at the end of the test? Report your answer using 3 significant figures.
Answer #1

K = 0.693 / t1/2

= 0.693 / 4.90

= 0.1414 s-1

Ln ( A° / A ) = rate constant * time

= ln ( 30.0 / A ) = 0.1414 * 3.00

A = 19.63 mg

Or 19.6 mg

If you have any questions please comment